CRY+VIS Interface Meeting
Date: 2022/3/4 13:00-13:30
Participants: Takahashi, Akutsu, Ikeda, Ushiba, Aso, Node, Sato, Miyoki, Hirata
- IX: Glitch in the Oplev signal was reduced in the circuit.
- EX: Health check.
- IY: Investigation of the suspension behavior.
- EY: Preparation for the accelerometers. The pump was set on 3/4.
- Replacement of flip cables: Stock of the feedthrough flange.
- TM height setting: ITMY.
- Oplev: Setup of the Oplev in EY was almost done. Not yet in IY.
- ACC: Installation was postponed in EY.
- Commissioning in air: Alignment of ITMY.
- Commissioning in vacuum: Lock time is about 30min.
- Overall commissioning:
- Acceptance check:
- PF actuator: At the in-vacuum commissioning.
- New vacuum pumps for V-chambers: Implementation in this week.
- Heater control: Test in IXV/EXV.
- Power source for Oplev: Check of the electric ports in 3/2-3.
- TMS: Installation of the balancer in 2/28-3/1.
Date: 2022/3/4 14:00-14:50
Participants: Takahashi, Ishizaki, Fabian, Ushiba, Tanaka, Aso, Hirata, Sato
- Investigating the suspension behavior in IY.
Bending test of the SiC pipe (Report).
For the BS:
- We inspected visually the BS payload and the BF in order to make sure no part of the suspension was too close to the security structure. It was all clear. Actually, the most of the security structure stop screws were too far away from the suspension. We brought closer a few of them hut others remain too far. We should move them to a more suitable place in the near future.
- The cable of the F0 fishing rod had a problem that we fixed. Probably, the origin of the problem was a loose screw in a connector.
- I gave information to Miyo-kun for the transfer function auto-measurement system.
In PRM, we tested the prototype of the cable clamps for the uppermost stage. The assembly works well but modifications are necessary:
- More places for cables are needed. We need at least nine and the prototype has only five.
- We need to devise a way to keep the cable shield, that is exposed, from touching the clamp base, which is made of aluminium.
Hirata-san and I will talk about these changes next week.
Other: we installed in ASC0 rack the QPD driver for the POP infrared transmitted beam.
Type-B and Type-Bp TO DO list
- Agree on the changes needed for the Type-Bp cable prototype.
- Provide more information to Miyo-kun for the auto-measurement system.
- Type-A team: Investigation of GAS filters in IY. Installation of ACCs in EY.
- Type-B team:
(Date in Kamioka)
The week from 3/7
- Takahashi: 3/7-11
- Sato: 3/7-11
- Hirata: 3/8-11
The week from 3/14
- Takahashi: 3/15-18
- Sato: 3/15-18
- Hirata: 3/15-18